OpenStudy (anonymous):

Translate and solve the algebraic statement for the inequality described below: “Six less than three times a number (n) is greater than or equal to 12.” n ≥ 6 n ≥ 4 n ≥ 2 n ≥ 1 @Joel_the_boss @cj49 @iGreen

JoelTheBoss (joel_the_boss):

I think it might be the first one

OpenStudy (cwrw238):

3n - 6 >= 12

OpenStudy (cwrw238):

to solve add 6 to both sides then divide both sides by 3

OpenStudy (igreen):

\(3n - 6 \ge 12\) Add 6 to both sides: \(3n \ge 18\) Divide both sides by 3: \(n \ge 6\)

OpenStudy (anonymous):

thanks

OpenStudy (igreen):

No problem.